# Cooperative Coevolution for Non-Separable Large-Scale Black-Box Optimization: Convergence Analyses and Distributed Accelerations

### v3

@@ Abstract
journal: arXiv
@@ 1 Introduction
Problem Structure: It is natural to deduce that not all non-separable problems can be handled efficiently by CC. For example, on many ill-conditioned non-separable landscapes, both CC and its gradient-based counterpart (i.e., coordinate descent, CD [59]) are typically worse than second-order-type optimizers (e.g., LM-CMA [60] and L-BFGS [61]). On the contrary, on some other non-separable landscapes (e.g., with relatively sparse variable interactions [62]), some modern CC could obtain very competitive (sometimes even state-of-the-art) results. Therefore, a key theoretical question arises: On what kinds of problem types from (non-separable) real-world problems CC is preferred over others? As argued in 1989 by Conn et al. [63], PAS functions “are clearly a very restricted case of partially separable (PS) functions [^5] We notice that there are several CC papers do not distinguish PAS and PS clearly, which may cause confusions. In the mathematical optimization (MO) community, PS includes not only PAS, but also non-separable functions with a sparse linkage structure. In effect, the focus of MO is the latter rather than the former, as the former is seen as “a very restricted case”, which is not emphasized by some CC papers.. Here we exclude PAS because of its hard-to-satisfy assumption in practice.
To the best of our knowledge, there have no theoretical work to satisfactorily solve the above theoretical challenge. In this paper, we first answer a related but much simpler practical question: When may end-users prefer to use divide-and conquer methods in practice? Some common answers are presented in the following:
@@
    There may be a relatively weak interaction between subcomponents in some (not all) non-separable real-world optimization problems, where decomposition-based methods could converge fast (within a practically accepted accuracy). This may partly explain why there always have some researchers and practitioners to use them since the establishment of the optimization area (furthermore, another advantage is their relative ease to understand and implement) [59, 71].
Convergence Analyses: In order to make it mathematically tractable, we simplify CC as a continuous-game (CG) model [72] but without losing its essential property. As compared to previous evolutionary game theory (EGT) [65, 73], our CG model can provide a much clearer analytical perspective built on only the pure Nash equilibrium (PNE) concept, much simpler than its mixed counterparts used in EGT. We theoretically show that under what conditions CC convergences, which most convex-quadratic benchmark functions can satisfy (see Section 3). To validate its prediction ability, we further demonstrate that many overlapping functions from the two newest test suites could essentially satisfy these conditions, which illustrates that CC finally converges to the global optimum on them, which is proved for the first time (our simulation experiments can perfectly match our predictions). Furthermore, using this new model, we reconfirm previously discovered pathologies of CC in a unified manner (i.e., relative generalization [74] and loss of gradients [75]). Depending upon these theoretical analyses, we propose to use a hierarchical decomposition structure to alleviate these issues and to obtain better generalizability (see Section 4.A).
@@ 3.1.2 Convergence to a PNE in CG
For successive transformation processes, the key point is to extract a strictly decreasing fitness subsequence from the nonincreasing fitness sequence of the best-so-far solution $\mathbf{x}$ generated by $\mathrm{CC}$. This could be guaranteed by a strong assumption that each suboptimizer has sufficient search ability for each subproblem (that is, it can find the global minimizer for each subproblem given a finite number of iterations). It is worthwhile noting that even such a strong assumption cannot ensure the convergence to the global optimum on the original problem (see Fig. 1 for an example). As a result, finally $\lim_{t\rightarrow+\infty}T_{t}(\mathbf{x})$ will move towards a fixed point (that is, ${T}(\mathbf{x})=\mathbf{x})$, which is also a PNE [^10] Its mathematical proof is simple via proof by contradiction: In fact, the Nobel-prize winner Nash was the first to connect the fixed point with the equilibrium point (now named after him) in his seminar PNAS and Annals of Mathematics papers, laying the theoretical foundation of modern game theory.. Please refer to A for a detailed mathematical demonstration. In practice, the global minimizer of each subproblem in each cycle is not needed to find exactly [62] in each cycle.
@@ 4 Distributed Multi-level Learning
In this section, a distributed cooperative coevolution (DCC) framework is proposed where the recent multilevel learning [77, 76] is employed, in order to improve the efficiency and effectiveness of CC on modern clustering computing platforms. Given the high complexity of distributed algorithms, we provide its source code available at GitHub [^14] https://github.com/Evolutionary-Intelligence/DCC to ensure repeatability. For distributed algorithms, multiple performance trade-offs must be considered: such as, distributed scheduling, distributed (shared) memory management, network communications, fault tolerant, system control, and so on. In this paper, we focus on the application-level parallelism and leave other tedious tasks to the underlying distributed computing engine.
@@ 4.1 Hierarchical Structure of DCC
As previously shown, CC tends to convergence to the PNE. If the PNE is not the global optimum (refer to Fig. 1 as an example), CC easily gets trapped into this suboptimal solution. This issue (known as relative generalization) is attributed to decomposition, which cannot be avoided in essence but may be alleviated somewhat. On a class of ill-conditioned non-separable functions [79], like its gradient-based counterpart (CD), CC often shows much slower convergence rates than second-order-type optimizers (e.g., LM-CMA and L-BFGS). However, when as a general-purpose BBO framework for LSO, it is highly desirable that CC could also handle these challenging issues. To achieve this goal, we use the recently proposed multilevel learning framework [77, 76] to combine the best of both worlds (i.e., the powerful invariance property of LM-CMA and the fine-tuning ability of CC via CMA-ES on low-dimensional subspace). In this paper, we only consider the hierarchical structure to implement it for simplicity, as seen in Fig. 3. In principle, a more general recursive structure can be also applied here, which we leave for a future work.
@@
As shown in Fig. 3, there are multiple slave nodes as well as one master node for distributed computing (e.g., on Spark [135] and Ray [136]). For DCC, at the meta-level the master node mainly coordinates different optimizers, each of which is run in one separate computing unit (a CPU core). Here we choose two different ES versions (CMA-ES [27] and LM-CMA [60]) as two search engines for the low-dimensional (often $<50$) subspace (after decomposition) and the original large-scale ($\gg 100$) space, respectively. In principle, any other EAs (e.g., GA [137], EP [138], PSO [139], DE [140], and EDA [141]) can be also selected as the suboptimizer. The rationale behind using two search engines lies that in the original space LM-CMA has a high possibility to escape from the suboptimal PNE via learning of promising evolution paths while maintaining a low computationally complexity (see [60] for analysis) and in the decomposed subspace the full-fledged CMA-ES [27] can well approximate the inverse of its Hessian matrix for fine-tuning with a reasonable memory consumption on each single computing unit (which is vital for scalability of distributed algorithms for LSO). Refer to the following three subsections as well as Algorithm 1 for more details.
@@ 4.2 Collective Learning of Covariance Matrices
where $\mathbf{d}^{t}\in\mathbb{R}^{n},\mathbf{I}\in\mathbb{R}^{n*n},\mathbf{p}^{t}\in\mathbb{R}^{n},\mathbf{z}^{t}\in\mathbb{R}^{n}\sim N(\mathbf{0},\mathbf{I}),0<c_{1}$ $<1$ are the directional vector, identity matrix, evolution path, standard Gaussian permutation, and learning rate of the rank-one update for iteration $t$, respectively. In the actual code implementation, $\mathbf{I}$ is omitted since $\mathbf{I}\mathbf{z}^{t}=\mathbf{z}^{t}$ The above form should be implemented from right to left, in order to avoid the too expensive matrix operation. Only $m=O(\log(n))\ll n$ dot products (with $O(n*\log(n))$ computational complexity) are involved when sampling each offspring.
@@ 5.2 High-Dimensional Test Functions
Since the focus of this paper is non-separable large-scale BBO, we choose 10 high-dimensional test functions, which are often used to analyze the convergence property in the ES community. For modern ESs, invariance is one fundamental design principle when optimizing non-separable functions. See Table 1 for their detailed formula. As a standard benchmarking [145] practice, we use the rotation and shift operation to make the test function non-separable and avoid the origin being the global optimum, respectively.
@@ 5.3 Experimental Setups
On all test functions, each algorithm was totally run 5 times. We set two termination conditions for all algorithms to make fair comparisons as much as possible: 1) when the best-so-far fitness is below $1\mathrm{e}{-10}$; 2) when the maximum runtime exceeds 3 hours. As a result, the total CPU runtime roughly equals 6600 (= 44 algorithms*10 functions*5 trails*3 hours) hours (i.e., 275 days) if they are run only on a single machine. To reduce the time-consuming benchmarking process, 10 HPC servers were used in parallel, all of which were also together used to build a (slightly heterogeneous) clustering computing platform. There is a total of 400 CPU logic cores and about 340GB memory in our private clustering computing platform. Under page limits, please see online Supplementary Materials (available at https://github.com/Evolutionary-Intelligence/DCC) for detailed hardware and software configurations.
@@ 5.4 Experimental Results and Analyses
Figure 4: Median convergence curves on a set of 2000-d rotated and shifted test functions. To avoid crowd and confusion (given 44 algorithms), only the first three of top-ranked optimizers have a legend and only half (22) of optimizers are drawn in each subfigure. Refer to online Supplementary Materials for detailed data.
Figure 5: Median convergence curves on a set of 2000-d rotated and shifted test functions. To avoid crowd and confusion (given 44 algorithms), only the first three of top-ranked optimizers have a legend and only half (22) of optimizers are drawn in each subfigure. Refer to online Supplementary Materials for detailed data.
In Fig. 4 and 5, the convergence curves on all test functions are drawn for 22 black-box optimizers. To avoid crowd in plotting, all other 22 optimizers are shown in online Supplementary Materials (available at https://github.com/Evolutionary-Intelligence/DCC), since all of them are much worse than or close (only on one test function) to our distributed algorithm on these high-dimensional functions.

### v4

@@ Abstract
journal: Applied Soft Computing
@@ 1 Introduction
Problem Structure: It is natural to deduce that not all non-separable problems can be handled efficiently by CC. For example, on many ill-conditioned non-separable landscapes, both CC and its gradient-based counterpart (i.e., coordinate descent, CD [59]) are typically worse than second-order-type optimizers (e.g., LM-CMA [60] and L-BFGS [61]). On the contrary, on some other non-separable landscapes (e.g., with relatively sparse variable interactions [62]), some modern CC could obtain very competitive (sometimes even state-of-the-art) results. Therefore, a key theoretical question arises: On what kinds of problem types from (non-separable) real-world problems CC is preferred over others? As argued in 1989 by Conn et al. [63], PAS functions “are clearly a very restricted case of partially separable (PS) functions [^5] We notice that there are several CC papers do not distinguish PAS and PS clearly, which may cause confusions. In the mathematical optimization (MO) community, PS includes not only PAS, but also non-separable functions with a sparse linkage structure. In effect, the focus of MO is on the latter rather than the former, as the former is seen as “a very restricted case”, which is not emphasized in some CC papers.. Here we exclude PAS because of its hard-to-satisfy assumption in practice.
To the best of our knowledge, there has no theoretical work to satisfactorily solve the above theoretical challenge. In this paper, we first answer a related but much simpler practical question: When may end-users prefer to use divide-and conquer methods in practice? Some common answers are presented in the following:
@@
    There may be a relatively weak interaction between subcomponents in some (not all) non-separable real-world optimization problems, where decomposition-based methods could converge fast (within a practically accepted accuracy). This may partly explain why there always have some researchers and practitioners using them since the establishment of the optimization area (furthermore, another advantage is their relative ease to understand and implement) [59, 71].
Convergence Analyses: In order to make it mathematically tractable, we simplify CC as a continuous-game (CG) model [72] but without losing its essential property. As compared to previous evolutionary game theory (EGT) [65, 73], our CG model can provide a much clearer analytical perspective built on only the pure Nash equilibrium (PNE) concept, much simpler than its mixed counterparts used in EGT. We theoretically show that under what conditions CC converges, which most convex-quadratic benchmark functions can satisfy (see Section 3). To validate its prediction ability, we further demonstrate that many overlapping functions from the two newest test suites could essentially satisfy these conditions, which illustrates that CC finally converges to the global optimum on them, which is proved for the first time (our simulation experiments can perfectly match our predictions). Furthermore, using this new model, we reconfirm previously discovered pathologies of CC in a unified manner (i.e., relative generalization [74] and loss of gradients [75]). Depending upon these theoretical analyses, we propose to use a hierarchical decomposition structure to alleviate these issues and to obtain better generalizability (see Section 4.A).
@@ 3.1.1 Ubiquity of Pure Nash Equilibrium (PNE)
In a nutshell, all global optima are PNE; however, the inverse does not necessarily hold (as in Fig. 1, we give a counterexample to prove it). Unfortunately, there is no certain mapping relationship between PNE and local optima: Some local optima may be PNE but some may be not; similarity some PNE may be local optima while some may be not (please refer to Fig. 1 for a counterexample). Here we choose to omit their rigorous proofs owing to their mathematical ease to design counterexamples. Note that since all local optima are global optima for convex functions, this greatly simplifies the global convergence analyses of CC on this specific yet important function form (see the Main Theorem for proof details).
@@ 3.1.2 Convergence to a PNE in CG
For successive transformation processes, the key point is to extract a strictly decreasing fitness subsequence from the nonincreasing fitness sequence of the best-so-far solution $\mathbf{x}$ generated by $\mathrm{CC}$. This could be guaranteed by a strong assumption that each suboptimizer has sufficient search ability for each subproblem (that is, it can find the global minimizer for each subproblem given a finite number of iterations). It is worthwhile noting that even such a strong assumption cannot ensure the convergence to the global optimum on the original problem (see Fig. 1 for an example). As a result, finally $\lim_{t\rightarrow+\infty}T_{t}(\mathbf{x})$ will move towards a fixed point (that is, ${T}(\mathbf{x})=\mathbf{x})$, which is also a PNE [^10] Its mathematical proof is simple via proof by contradiction: In fact, the Nobel-prize winner Nash was the first to connect the fixed point with the equilibrium point (now named after him) in his seminar PNAS and Annals of Mathematics papers, laying the theoretical foundation of modern game theory.. Please refer to A for a detailed mathematical demonstration. In practice, the global minimizer of each subproblem in inner cycle is not needed to find exactly [62] in outer cycle.
@@ 4 Distributed Multi-level Learning
In this section, a distributed cooperative coevolution (DCC) framework is proposed where the recent multilevel learning model for biological evolution [77, 76] is employed, in order to improve the efficiency and effectiveness of CC on modern clustering computing platforms. Given the high complexity of distributed algorithms, we provide its source code available at GitHub [^14] https://github.com/Evolutionary-Intelligence/DCC for our proposed framework to ensure repeatability. For distributed algorithms, multiple performance trade-offs must be considered: such as, distributed scheduling, distributed (shared) memory management, network communications, fault tolerant, system control, and so on. In this paper, we focus on the application-level parallelism and leave other tedious tasks to the underlying distributed computing engine.
@@ 4.1 Hierarchical Structure of DCC
As previously shown, CC tends to convergence to the PNE. If the PNE is not the global optimum (refer to Fig. 1 as an example), CC easily gets trapped into this suboptimal solution. This issue (known as relative generalization) is attributed to decomposition, which cannot be avoided in essence but may be alleviated somewhat. On a class of ill-conditioned non-separable functions [79], like its gradient-based counterpart (CD), CC often shows much slower convergence rates than second-order-type optimizers (e.g., LM-CMA and L-BFGS). However, when as a general-purpose BBO framework for LSO, it is highly desirable that CC could also handle these challenging issues. To achieve this goal, we use the recently proposed multilevel learning framework [77, 76] to combine the best of both worlds (i.e., the powerful invariance property of LM-CMA and the fine-tuning ability of CC via CMA-ES on multiple low-dimensional subspace). In this paper, we only consider using the hierarchical structure to implement it for simplicity, as seen in Fig. 3. In principle, a more general recursive structure can be also applied here, which we leave for a future work.
@@
As shown in Fig. 3, there are multiple slave nodes as well as one master node for distributed computing (e.g., on Spark [135] and Ray [136]). For DCC, at the meta-level the master node mainly coordinates different optimizers, each of which is run in one separate computing unit (i.e., a single CPU core). Here we choose two different ES versions (CMA-ES [27] and LM-CMA [60]) as two search engines for the low-dimensional (often $<50$) subspace (after decomposition) and the original large-scale ($\gg 100$) space, respectively. In principle, any other EAs (e.g., GA [137], EP [138], PSO [139], DE [140], and EDA [141]) can be also selected as the suboptimizer. The rationale behind using two search engines lies that in the original space LM-CMA has a high possibility to escape from the suboptimal PNE via learning of promising evolution paths while maintaining a low computationally complexity (see [60] for analysis) and in the decomposed subspace the full-fledged CMA-ES [27] can well approximate the inverse of its Hessian matrix for fine-tuning with a reasonable memory consumption on each computing unit (which is vital for scalability of distributed algorithms for LSO). Refer to the following three subsections as well as Algorithm 1 for more details.
@@ 4.2 Collective Learning of Covariance Matrices
where $\mathbf{d}^{t}\in\mathbb{R}^{n},\mathbf{I}\in\mathbb{R}^{n*n},\mathbf{p}^{t}\in\mathbb{R}^{n},\mathbf{z}^{t}\in\mathbb{R}^{n}\sim N(\mathbf{0},\mathbf{I}),0<c_{1}$ $<1$ are the directional vector, identity matrix, evolution path, standard Gaussian permutation, and learning rate of the rank-one update for iteration $t$, respectively. In the actual code implementation, $\mathbf{I}$ is omitted since $\mathbf{I}\mathbf{z}^{t}=\mathbf{z}^{t}$ The above form should be implemented from right to left, in order to avoid the computationally expensive matrix operation. Only $m=O(\log(n))\ll n$ dot products (with $O(n*\log(n))$ computational complexity) are involved when sampling each offspring.
@@ 4.4 Distribution Mean Update via Weighted Averaging
4.5 Convergences and Implementations
Since our proposed hierarchical CC framework uses the best-so-far solution as the fitness evaluation base of subpopulations as the same as the continuous-game model, its convergence curve can be also theoretically modeled under the Main Theorem in Section 3.2 without significant modifications. As a result, the global convergence condition of the Main Theorem can be well transferred to our new hierarchical CC framework directly.
On ill-conditioned non-separable landscapes, CC suffers easily from the much slow rate of convergence owing to strong dependencies among variables. However, on many non-separable landscapes with weak and sparse variable interactions, some modern CC could obtain very competitive and sometimes even state-of-the-art optimization performances. This is because weak and sparse interactions between subcomponents (aka well-conditioned landscapes) could approximately match the underlying assumptions behind CC. When as a general-purpose optimizer for large-scale BBO, it is highly desirable for CC to handle different fitness landscapes as many as possible. For this important aim, our proposed hierarchical CC framework combines the fine-tuning ability from CC’s decomposition (on well-conditioned landscapes) with the topology learning (aka invariance) ability of LM-CMA (in particular on ill-conditioned landscapes), in order to enjoy the best of worlds. For local optima avoidance, restart is often a simple yet efficient strategy, which can be easily integrated into our CC framework.
@@ 5.2 High-Dimensional Test Functions
Since the focus of this paper is non-separable large-scale BBO, we choose 10 high-dimensional test functions, which are often used to analyze the convergence property in the ES community. For modern ESs, invariance is one of fundamental design principles when optimizing non-separable functions. See Table 1 for their detailed formula. As a standard benchmarking [145] practice, we use the rotation and shift operation to make all test functions non-separable and avoid the origin being the global optimum, respectively.
@@
For fair comparisons of all 42 general-purpose black-box optimizers, we did not fine-tune any hyper-parameter settings for each optimizer on every test function independently. Instead, arguably as a standard practice, we used the same hyper-parameter values on all test functions. For all 41 baseline optimizers, we used the default values suggested from the well-designed open-source library called PyPop7, since at least these default values have been well-tested in their corresponding papers. For our own distributed algorithm, the maximum of available parallel computing units (independent CPU cores considered in our paper) is a key factor to impact the proper settings of its hyper-parameters. For simplicity, it is better to self-adapt these hyper-parameters according to the maximum of available parallel computing resources, if possible. Specifically, the total number of parallel sub-optimizers is set as the total number of available CPU (logical) cores minus the number of their corresponding (Linux) servers, in order to avoid possible program blocking owing to competition of computing resources. This is mainly because we need to leave (at least) one CPU core in each independent (Linux) server to do a series of necessary OS-level management tasks such as network communications, task scheduling, and so on.
@@ 5.3 Experimental Setups
On all test functions, each algorithm was totally run 5 times. We set two termination conditions for all algorithms to make fair comparisons as much as possible: 1) when the best-so-far fitness is below $1\mathrm{e}{-10}$; 2) when the maximum runtime exceeds 3 hours. As a result, the total CPU runtime roughly equals 6450 (= 43 algorithms*10 functions*5 trails*3 hours) hours (i.e., 269 days) if they were run only on a single machine. To reduce the time-consuming benchmarking process, 10 HPC servers were used in parallel, all of which were also together used to build a (slightly heterogeneous) clustering computing platform. There are a total of 400 CPU logic cores and about 340GB memory in our private clustering computing platform. Under page limits, please see online Supplementary Materials (available at https://github.com/Evolutionary-Intelligence/DCC) for detailed hardware and software configurations.
@@ 5.4 Experimental Results and Analyses
Figure 4: Median convergence curves on a set of 2000-d rotated and shifted test functions. To avoid crowd and confusion (given 44 algorithms), only the first half (22) of optimizers are drawn in each subfigure. Refer to online Supplementary Materials for detailed data.
In Fig. 4, the convergence curves on all test functions are drawn for 22 black-box optimizers. To avoid crowd in plotting, all other 22 optimizers are shown in online Supplementary Materials (available at https://github.com/Evolutionary-Intelligence/DCC), since all of them are much worse than or close (only on one test function) to our distributed algorithm on these high-dimensional functions.
@@
6 One Application Case on Black-Box Classification
To validate the practical utility of our distributed framework, in this section we choose one application case from data science as the testbed, namely black-box classification. Currently, (two-class) black-box classification is commonly used to compare the performance of different zeroth-order optimizers in more realistic scenarios. Here our used dataset called QSAR androgen receptor (QAR) comes from the biological chemistry community, which is accessible at the popular UCI Machine Learning Repository. For this dataset, there are a set of 1687 instances with 1024 features.
Here we adopt a common loss function
$g(\mathbf{w})=\sum_{p=1}^{P}\max\left(0,1-y_{p}\stackrel{{\scriptstyle\circ}}{{\mathbf{x}}}^{T}\mathbf{w}\right)$
where $\left\{\left(\mathbf{x}_{p},y_{p}\right)\right\}_{p=1}^{P}$ with the labels $y_{p}\in\{-1,+1\}$ represents training dataset, $\stackrel{{\scriptstyle\circ}}{{\mathbf{x}}}^{T}\mathbf{w}=0$ is a linear decision boundary, from support vector machines (SVM). Note that there exist non-differentiable search points, since this loss function involves a non-smoothed max() operation. Therefore, this loss function has been commonly used for black-box classification. For the dataset QAR, there are 1025 (= 1024 + 1) weights to be optimized, after we consider the bias term in the SVM loss function.
Figure 5: SVM loss function
As is shown in Fig. 5, our distributed algorithm (DCC) can obtain the fastest rate of convergence and the best quality of solution on this high-dimensional loss function, among 9 competitive algorithms. The reason behind DCC’s superiority on this application case may lie in that our distributed algorithm could efficiently utilize these parallel computing units to collectively learn its topology/geometry structure well. In the future, we plan to test more application cases to validate the potential of DCC.
@@ 7 Conclusion
Furthermore, formalizing CC as a continuous game model via simplifications suffers from two main issues: 1) it does not consider other complex subcomponent interactions (e.g., lenient learners) except the widely used interaction strategy (i.e., based on only the best-so-far solution); 2) it relies heavily on an idealized assumption upon all sub-optimizers, that is, each sub-optimizer could find the optima for all subproblems under a limited (even possibly a very huge) number of function evaluations. We leave the relaxation of these two theoretical assumptions for further works.